# A Novel Passive Occupational Shoulder Exoskeleton With Adjustable Peak Assistive Torque Angle For Overhead Tasks

### v1

@@ 1 Introduction
OSEs can be classified into active, semi-active, and passive categories. In general, active OSEs can provide higher, more precise, and versatile assistive torques [8, 9, 10]. However, they have larger volumes and higher maintenance costs [11]. As an alternative solution, semi-active OSEs can automatically adjust torque profiles by employing small low-power motors [12], such as the H-PULSE [9], [13]. Compared to active ones, it is more compact and has lower energy consumption while maintaining effectiveness. However, compared to passive ones, the maintenance and manufacturing costs of semi-active OSEs are higher, which would limit their widespread adoption and application. Passive OSEs (POSEs) are characterized by their low cost and lightweight design [14], [15], which provide arm gravity compensation during overhead works without external power sources. Recent studies have demonstrated that wearing POSEs can reduce muscle activation during overhead tasks performed in laboratory and field settings [16, 17, 18]. Even though POSEs have been commercialized, some pending issues remain, thus impeding user acceptance and assistive efficiency [19].
@@
One of the main issues is the insufficient kinematic compatibility between the exoskeleton and the user [20]. POSEs can be classified based on the presence or absence of an exoskeletal shoulder structure above the human shoulder [11]. Exoskeletons with a rigid structure above the human shoulder, such as the ShoulderX [21] and Airframe [22], typically feature a degree of freedom (DoF) next to the shoulder to follow shoulder flexion/extension (F/E) movements. These exoskeletons are compact and thus are suitable for performing overhead tasks in narrow spaces. However, they may induce collisions between the shoulder and the exoskeleton at high elevation angles, resulting in an insufficient range of motion (ROM). Exoskeletons without a structure above the shoulder, such as the SPM [23] and Exo4Work [24], avoid collisions at high elevation angles.However, they generally incorporate multi-DoF structures in the back to accommodate shoulder movements, which leads to a larger volume. The drawback may result in collisions with the surrounding objects in narrow spaces. In existing exoskeletons, achieving both sufficient ROM and compactness is mutually exclusive. To achieve a balance between them, Junsoo Kim et al. [11] designed an exoskeleton with tilted shoulder structures for assisting overhead tasks. However, it still cannot provide a sufficient ROM for horizontal F/E movements, which may induce a decreased user experience. There is an urgent need for an exoskeleton that can simultaneously provide sufficient ROM and compactness.
Figure 1: The torque profiles when there is a match (orange line) or mismatch (blue line) between the PATA and the target angle of the task in an overhead task with a target angle of 120º. Angle difference refers to the discrepancy between the PATA and target angle of the task.
@@
Another main issue is the task-specific or individual-specific design, corresponding to a fixed assistive torque profile and thus a fixed assistive region (i.e. the subset of ROM where the exoskeleton can assist), which may not be optimal for other tasks [25], [26]. Overhead tasks relate to different positions, especially different elevation angles of the shoulder (e.g., screwing task of the automobile manufacturing industry [21, 27, 28], aircraft squeeze riveting and sealing tasks [29], [30], the welding task of shipbuilding [11], order picking task [16] and so on). Consequently, various overhead tasks require different optimal torque profiles [13], [24]. In addition, the same task may have different target angles for individuals of different heights, which requires providing personalized torque profiles [31]. The torque profile should be adaptive to tasks and subjects, essentially characterized by the amplitude and peak assistive torque angle (PATA). The amplitude of the peak assistive torque can be easily tuned by adjusting the extension length of the spring. For example, the MATE can manually change the peak torque by a knob [32], and H-PULSE can adjust the spring length by a spindle drive coupled with a servomotor [9], [13]. The timing of the peak assistive torque ensures that the POSE can provide timely assistance when the torque profile reaches its peak value at specific elevation angles of the shoulder [19]. Fig. 1 depicts the torque profiles when there is a match or mismatch between the PATA and the target angle of the task in an overhead task with a target angle of 120º. When there is a match between them, it ensures timely assistance and reduces energy loss, making it a more efficient assistive strategy. However, current POSEs can only change the PATA through reconstruction, which means they cannot provide torque profiles for different tasks or individuals, i.e., insufficient adaptability [26], [31]. In summary, different tasks and individuals should have different PATAs, lacking in current POSEs. Therefore, an adjustable PATA is needed to enhance task and subject adaptability.
@@
The object of this study is to develop, implement, and preliminarily evaluate a novel POSE for overhead tasks, named HIT-POSE. Our prototype can provide a sufficient ROM of the human shoulder and possess an adjustable PATA, thus adapting to various overhead tasks and individuals.The major contributions of our work are as follows:
@@ 2.1 System Overview
Overhead work is defined as working with hands above acromion or over 60° shoulder flexion or abduction [33]. Therefore, we can define the workspace for overhead tasks as the sagittal flexion of the shoulder with 60° to 168.8° (the maximum shoulder flexion angle for healthy adult male [34]), abduction angle of the shoulder greater than 60°, and horizontal flexion of the shoulder with 0° to 162.2° (the maximum shoulder horizontal flexion angle for healthy adult male [11]). In addition, the total weight of the exoskeleton should be less than 5kg [19].
HIT-POSE is designed to compensate for shoulder gravitational torque while providing sufficient ROM for the shoulder, as illustrated in Fig. 2. The total weight of the device 2.8 kg. It is suitable for people with a back length of 43.6-66.6 cm and a shoulder width of 29.6-45.6 cm, and the volume of our device is 22cm × 48cm × 15cm. HIT-POSE comprises four main modules: (a) the ergonomic shoulder structure, (b) the torque generator with adjustable PATA, (c) the physical user-exoskeleton interface (PUEI), and (d) the size regulation module.
@@ 2.2 Ergonomic Shoulder Structure
We conducted kinematic modeling of the exoskeleton and the user and simulated ROM to select optimal design parameters. The proposed exoskeleton has a horizontal F/E joint above the shoulder and lacks redundant structures in the back. Our object here is to provide sufficient ROM for the shoulder, especially at high elevation angles.
The intersection of the rotational axes of the sagittal and horizontal F/E joints is considered the center of the shoulder of the exoskeleton (CSE), as shown in Fig. 3. The CSE is close to the center of the shoulder of the user (CSU), enabling the torque generator to trace the shoulder movements with a relatively low DoF (Fig. 3). However, it does not coincide with the CSU, indicating the necessity to find the optimal design parameters. As explained in [35] and [36], the shoulder motion can be observed with humeral elevation and posterior displacement of the CSU.
@@ 2.3 Torque Generator with Adjustable PATA
The torque generator is composed of a sagittal F/E joint bar, PATA adjustable module, parallel spring group, spring pretension tuning module, carbon fiber shell, and five pulleys with inner bearings (Fig. 6a). A wire rope of 1mm diameter connects the parallel spring group to pulley 2. The pulleys 3 and 4 can be moved vertically through the guide rail, resulting in the adjustable PATA. The 8-shaped loop with a rotational DoF connects the parallel spring group with the spring pretension tuning module. It ensures that the parallel spring group does not experience axial rotation when screwing the spring pretension tuning module. Three identical springs (i.e., $L_{s}$=$L_{AB}$=$L_{CD}$=$L_{EF}$) are installed in the parallel spring group to obtain a high energy storage capability.
@@
Based on (1) to (3) and Fig. 7b, the main factor affecting PATA is $L_{exo}$, rather than $\Delta{L}$, as it is monotonically decreasing. The primary parameters affecting the phase shift of $L_{exo}$ are $\alpha$ and $\beta$, denoting the angles of pulleys 2 and 3, as shown in Fig. 6b and 6c. When $\alpha$ decreases or $\beta$ increases, PATA would decrease and thus the phase of $L_{exo}$ curve would shift to the left (Fig. 7a and 7b). The assistive torque would be influenced by $L_{exo}$ and also have a phase shift (as shown in Fig.1). To better investigate the trend of PATA, when $\alpha$=$\beta$=0, we defined the PATA as a reference, denoted by $\theta_{0}$. The PATA ($\theta_{PATA}$) can be derived as:
@@ 2.4 Parameters Impact on Torque Profiles
As visualized in Fig. 6b and 6c, the generated assistive torque can be divided into two phases based on whether the wire rope is routed to the sagittal F/E joint bar. When the wire rope is attached to the sagittal F/E joint bar (Phase $\mathrm{I}$, Fig. 6b), the assistive torque remains low to avoid hindering normal activities like walking, corresponding to the low assistance zone (Fig. 7b). In contrast, when the wire rope is detached from the sagittal F/E joint bar (Phase $\mathrm{II}$, Fig. 6c), the assistive torque increases rapidly to provide effective assistance for the target task, corresponding to the high assistance zone (Fig. 7b). We defined the critical angle for switching between phases $\mathrm{I}$ and $\mathrm{II}$ as $\theta_{c}$ (Fig. 7a), which is dependent on $\alpha$ and $\beta$ (Fig. 6b and 6c). As $\alpha$ increases or $\beta$ decreases, $\theta_{c}$ increases (Fig. 9e - 9f). The wire rope coupled with the sagittal F/E joint bar formulates a passive clutch mechanism that allows for switching between low and high assistance zones. If there is no sagittal F/E joint bar (i.e., orange dotted line in Fig. 6b), the torque generator will pull the arm backward, which poses a safety hazard to the user. It is worth noting that this mechanism also provides a torque of approximately 2 N·m, although it is much smaller than the peak assist torque, it may still reduce the user’s acceptance of the exoskeleton.
@@ 2.5 PUEI and Size Regulation Module
The size regulation module is designed to accommodate diverse ergonomic needs while ensuring both comfort and safety. The height and width of the back are adjustable through upper and lower back linkages, as well as lateral back slides. Additionally, the flexible fabric allows for precise tightening of the exoskeleton around the user’s waist, back, and arms.
@@ 3.2 Data Recordings
ROM of the right-side shoulder was measured by the motion capture system (100Hz, VICON, Oxford Metrics, Oxford, UK). Muscular activities were monitored using the surface EMG electrodes ($1$Hz, Delsys Inc., Natick, MA, USA). Meanwhile, acceleration signals were recorded through an IMU (148Hz, Delsys Inc., Natick, MA, USA). Notably, the motion capture system did not work simultaneously with EMG. Thus, we did not synchronize them through a trigger.
@@ 3.3 Experimental Tasks
In session 1, participants were directed to maximize sagittal and horizontal F/E movements, at a self-selected constant pace, as shown in Fig. 10a. In session 2, participants were guided to screw the M6 bolts with a hex key (mass: 8.6 g), as shown in Fig. 10b. Nuts were embedded in an aluminum frame, which was set to two heights: low height was defined as hand with the shoulder at about 90º, and high height was defined as hand with the shoulder at about 120º (Fig. 11).
@@ 3.4 Procedure
Before the experiment, participants were introduced to the experimental paradigm and familiarized with HIT-POSE and tasks. In session 1, tasks were performed under two conditions: without the exoskeleton (NO HIT-POSE) and with the exoskeleton (HIT-POSE). Each participant completed six trials: three in each condition, with each trial consisting of five movements and lasting about 30 seconds. In session 2, the exoskeleton was configured for screwing at low and high heights (Fig. 12). Both configurations had the same peak torque but different PATAs for the respective heights. A match between the PATA and the task angle was defined as HIT-POSE Match; otherwise, it was HIT-POSE Mismatch. Each task was performed under three conditions: NO HIT-POSE, HIT-POSE Match, and HIT-POSE Mismatch (Fig. 12), with each trial lasting 2 minutes per condition.
Surface EMG electrodes were placed unilaterally on the right-side muscles according to SENIAM guidelines [39]: anterior deltoid (AD), middle deltoid (MD), posterior deltoid (PD), bicep brachii (BB), pectoralis major (PM), trapezius (TR), latissimus dorsi (LD), erector spinae (ES). Additionally, an IMU was placed on the right wrist to measure the acceleration. Then, the EMG signals were normalized for each participant and task using the maximal amplitude of each muscle during NO HIT-POSE conditions [40, 41, 42]. The order of conditions and tasks was randomized across participants to avoid order effects. After each trial, participants completed the NASA-TLX [43] to assess perceived workload (PW) and the Borg CR-10 [44] for rating perceived exertion (RPE). After all exoskeleton trials, they completed the System Usability Scale (SUS) [45]. Fig. 12 presents the experimental procedure for session 2.
@@ 3.5 Data and Statistical Analysis
In session 1, mean values of the maximum sagittal and horizontal F/E angles were calculated using shoulder angles measured by VICON and corresponding hand workspace were calculated using the position of the reflective point on hand measured by VICON among all participants in NO HIT-POSE and HIT-POSE conditions. In session 2, acceleration signals were resampled to $1$Hz, and variances were calculated using a 30ms sliding window. A set variance threshold determined the start and end times of each trial. EMG signals were processed to compute the linear envelopes by setting them to zero-mean, applying band-pass filter (4th, 20-350Hz), full-wave rectification, low-pass filter (6Hz), and normalization. For each trial, the root mean square (RMS) of each muscle was obtained from the linear envelope via a 30ms time sliding window during the phase of task execution.
Statistical analyses were used to assess the effectiveness of the exoskeleton in reducing muscle activations and subjective evaluation indicators (i.e., RPE and PW scores). The Shapiro-Wilk test was employed to check data normality. If the normal distribution was not satisfied, the non-parametric analysis of variance (ANOVA) was conducted to check for across-condition differences, otherwise, the parametric ANOVA was applied. A post-hoc analysis was conducted when a significant effect was observed. Statistical significance was concluded when p $<$ 0.05.

### v2

@@ 1 Introduction
OSEs can be classified into active and passive categories. In general, active OSEs can provide higher, more precise, and versatile assistive torques [8, 9, 10]. However, they have larger volumes, resulting in bulkier and heavier, and are prone to malfunctions (i.e., increased maintenance costs) and collision [11]. Passive OSEs (POSEs) characterize as cheaper, lightweight design and a trade-off between effectiveness and wearability [12] [13], thus are increasingly accepted by several manufacturers in real-application scenarios. They can provide arm gravity compensation during overhead works without external power sources. Recent studies have demonstrated that wearing POSEs can reduce muscle activation during overhead tasks performed in laboratory and field settings [14, 15, 16]. Overhead tasks relate to different positions, especially different elevation angles of shoulder (e.g., screwing task of the automobile manufacturing industry [17, 18, 19], aircraft squeeze riveting and sealing tasks [20, 21], the welding task of shipbuilding [22], order picking task [14] and so on). Even though POSEs have been commercialized, some pending issues still remain, thus impede the user acceptance and assistive efficiency [23].
@@
One of the main issues is the insufficient kinematic compatibility between the exoskeleton and the user [23]. POSEs can be classified into two types based on whether there is an exoskeletal shoulder structure above the human shoulder [11]. Exoskeletons with a rigid structure above the human shoulder, such as the ShoulderX [17] and Airframe [24], typically feature a degree of freedom (DoF) next to the shoulder to follow shoulder flexion/extension (F/E) movements. These exoskeletons are compact and thus are suitable for performing overhead tasks in narrow spaces. However, they may induce collisions between the shoulder and the exoskeleton at high elevation angles, resulting in an insufficient range of motion (ROM). Exoskeletons without a structure above the shoulder, such as the MATE [25], Exo4Work [26], and H-PULSE [9, 12], avoid collisions at high elevation angles. However, they generally incorporate multi-DoF structures in the back to accommodate shoulder movements, which leads to a larger volume. The drawback may result in collisions with the surrounding objects in narrow spaces. In existing exoskeletons, achieving both sufficient ROM and compactness is mutually exclusive. To achieve a balance between them, Junsoo Kim et al. [11] designed an exoskeleton with tilted shoulder structures for assisting overhead tasks. However, it still cannot provide a sufficient ROM for horizontal F/E movements, which may induce a decreased user experience. There is an urgent need for an exoskeleton that can simultaneously provide sufficient ROM and compactness.
@@
Another main issue of current POSEs is the task-specific or individual-specific design, corresponding to a fixed assistive torque profile and thus a fixed assistive region (i.e. the subset of ROM where the exoskeleton can provide assistance), which may not be optimal for other tasks [27, 22]. Considering that the assistive torque profile for overhead tasks are mainly inverse U-shaped due to the need of compensating for the gravity of arm elevation and lowering, the torque profile should be adaptive to tasks and subjects, essentially characterized by the amplitude of the peak torque and the timing of peak torque (i.e. the peak assistive torque angle, PATA). Different overhead tasks correspond to different postures and PATAs, thus require different optimal torque profiles [12, 25]. In addition, the same task may have different PATAs for individuals of different heights, which requires providing personalized torque profiles [28]. The amplitude of the peak assistive torque can be easily tuned by adjusting the extension length of the spring. For example, the MATE can manually change the peak torque by a knob [25], and the H-PULSE can adjust the spring length by a spindle drive coupled with a servomotor [9, 12]. The timing of the peak assistive torque ensures that the POSE can provide timely assistance when the shoulder torque reaches its peak value at some specific elevation angles of shoulder [23]. Fig. 1 depicts the torque profiles when the PATA matches or mismatches the requirement of the task in an overhead task with a target angle of 120º. When the PATA matches the target angle of the task, it ensures timely assistance and reduces energy loss, making it a more efficient assistive strategy. However, current POSEs can only change the PATA through reconstruction, which means they cannot provide torque profiles for different tasks or individuals, i.e., insufficient adaptability [22, 28]. In summary, different tasks and individuals should be matched with different PATA, lacking in current POSEs. Therefore, a POSE with adjustable PATA is needed to enhance its task and subject adaptability.
@@
Inspired by the issues mentioned above, the object of this study is to develop, implement, and preliminarily evaluate a novel POSE for overhead tasks, named HIT-POSE. Our prototype can provide sufficient ROM of the human shoulder and possess adjustable PATA, thus adapting to various overhead tasks and individuals. The major contributions of our work are as follows:
@@
The article is organized as follows. Section 2 introduces the structure design of HIT-POSE in detail. Section 3 presents the experimental evaluation methods. The evaluation results are reported in Section 4. Section 5 demonstrates the discussion of the results, followed by a conclusion of the full article.
Figure 1: The torque profiles when the PATA matches (orange line) or mismatches (blue line) the target angle of the task in an overhead task with a target angle of 120º.
@@ 2.1 System Overview
The HIT-POSE is designed to compensate for shoulder gravitational torque while providing sufficient ROM for the shoulder, as illustrated in Fig. 2. The total weight of the device is 2.8 kg. The HIT-POSE comprises four main modules: (a) the ergonomic shoulder structure, (b) the torque generator with adjustable PATA, (c) the physical user-exoskeleton interface (PUEI), and (d) the size regulation module. The guiding principles of design are safety, compactness, efficiency, adequate ROM, and wearability. HIT-POSE can adjust the PATA and the peak torque to generate more versatile and efficient torque profiles, which can effectively assist the shoulder in multiple overhead tasks with different levels of arm elevations.
@@ 2.2 Ergonomic Shoulder Structure
In this subsection, we conducted kinematic modeling of the exoskeleton and the user and simulated ROM to select optimal design parameters, and therefore ensure sufficient ROM for the shoulder when wearing the exoskeleton. The shoulder bending linkage is connected to the torque generator and front-and-rear adjustable linkage through the sagittal and horizontal F/E joints, respectively (as shown in Fig. 2). The proposed exoskeleton is a POSE with a horizontal F/E joint above the shoulder and without redundant structures in the back. Our object here is to provide sufficient ROM for the shoulder, especially at high elevation angles.
The intersection of the rotational axes of the sagittal and horizontal F/E joints is considered the center of the shoulder of the exoskeleton (CSE), as shown in Fig. 3. The CSE is close to the center of the shoulder of the user (CSU), enabling the torque generator to trace the shoulder movements with a relatively low DoF (Fig. 3). However, it does not coincide with the CSU, indicating the necessity to find the optimal design parameters. As explained in [29] and [30], the shoulder motion can be observed with humeral elevation and posterior displacement of the CSU. This phenomenon introduces great design inspiration to us.
@@ 2.3 Torque Generator with Adjustable PATA
The torque generator is composed of a sagittal F/E joint bar, PATA adjus module, parallel spring group, spring pretension tuning module, carbon fiber shell, and five pulleys with inner bearings (Fig. 6a). A wire rope of 1mm diameter connects the parallel spring group to pulley 2. The pulleys 3 and 4 can be moved vertically through the guide rail, resulting in the adjustable PATA. The 8-shaped loop with a rotational DoF connects the parallel spring group with the spring pretension tuning module. It ensures that the parallel spring group does not experience axial rotation when screwing the spring pretension tuning module. Three identical springs (i.e., $L_{s}$=$L_{AB}$=$L_{CD}$=$L_{EF}$) are installed in the parallel spring group to obtain a high energy storage capability.
@@
Based on (1) to (3) and Fig. 7b, the main factor affecting PATA is $L_{exo}$, rather than $\Delta{L}$, as it is monotonically decreasing. The primary parameters affecting the phase shift of $L_{exo}$ are $\alpha$ and $\beta$, denoting the angles of pulleys 2 and 3, as shown in Fig. 6b and 6c. As shown in Fig. 7 (a) and (b), when $\alpha$ decreases or $\beta$ increases, PATA would decrease and thus the phase of $L_{exo}$ curve would shift to the left. The assistive torque would be influenced by $L_{exo}$ and also have a phase shift (as shown in Fig.1). To better investigate the trend of PATA, when $\alpha$=$\beta$=0, we defined the PATA as a reference, denoted by $\theta_{0}$. The PATA ($\theta_{PATA}$) can be derived as:
@@ 2.4 Parameters Impact on Torque Profiles
As visualized in Fig. 6b and 6c, the generated assistive torque can be divided into two phases based on whether the wire rope is routed to the sagittal F/E joint bar. When the wire rope is attached to the sagittal F/E joint bar (Phase $\mathrm{I}$, Fig. 6b), the assistive torque remains low to avoid hindering normal activities like walking, corresponding to the low assistance zone (Fig. 7b). In contrast, when the wire rope is detached from the sagittal F/E joint bar (Phase $\mathrm{II}$, Fig. 6c), the assistive torque increases rapidly to provide effective assistance for the target task, corresponding to the high assistance zone (Fig. 7b). We defined the critical angle for switching between phases $\mathrm{I}$ and $\mathrm{II}$ as $\theta_{c}$ (Fig. 7a), which is dependent on $\alpha$ and $\beta$ (Fig. 6b and 6c). As $\alpha$ increases or $\beta$ decreases, $\theta_{c}$ increases (Fig. 9e - 9f). The wire rope coupled with the sagittal F/E joint bar formulates a passive clutch mechanism that allows for switching between low and high assistance zones.If there is no sagittal F/E joint bar (i.e., orange dotted line in Fig. 6b), the torque generator will generate torque to pull the arm backward, which poses a safety hazard to the user.
@@ 2.5 PUEI and Size Regulation Module
The size regulation module aims to meet different ergonomic requirements and guarantee comfort and safety. The height and width of the back can be controlled with the upper and lower back linkages and lateral back slides, respectively. The flexible fabric can be adjusted to tighten the exoskeleton to the waist, back, and arm of the user.
@@ 3.2 Data Recordings
ROM of the right-side shoulder was measured by the motion capture system (100Hz, VICON, Oxford Metrics, Oxford, UK) to verify whether the exoskeleton can provide sufficient ROM. Muscular activities were monitored using the surface EMG electrodes ($1$Hz, Delsys Inc., Natick, MA, USA) to explore the impact of the exoskeleton on biomechanical effort reduction. Meanwhile, acceleration signals were recorded through an IMU (148Hz, Delsys Inc., Natick, MA, USA). Notably, the motion capture system did not work simultaneously with EMG. Thus, we did not synchronize them through a trigger.
@@ 3.3 Experimental Tasks
To verify that HIT-POSE can provide sufficient ROM (session 1), participants were directed to maximize sagittal and horizontal F/E movements, at a self-selected constant pace, as shown in Fig. 10(a). For validating the performance of HIT-POSE (session 2), participants were guided to screw the M6 bolts with a hex key (mass: 8.6 g), as shown in Fig. 10(b). Nuts were embedded in an aluminum frame, which was set to two heights: low height was defined as hand with the shoulder at about 90º, and high height was defined as hand with the shoulder at about 120º (Fig. 11).
@@ 3.4 Procedure
Prior to the start of the experiment, participants were introduced to the experimental paradigm and guided to familiarize HIT-POSE and tasks. In session 1, tasks were performed under two conditions: without the exoskeleton (NO HIT-POSE) and with the exoskeleton (HIT-POSE). Each participant was required to complete 6 trials (3 trials were repeated in each condition). In session 2, the exoskeleton had two configurations for screwing at low and high heights, respectively (Fig. 12). The two configurations present the same peak torque but different PATAs corresponded to the heights. The condition where the PATA of the exoskeleton aligned with the PATA required for the height, was referred to as HIT-POSE Match. Otherwise, it was defined as HIT-POSE Mismatch. Each task was performed in 3 conditions: NO HIT-POSE, HIT-POSE Match, and HIT-POSE Mismatch (Fig. 12).
Surface EMG electrodes were placed unilaterally on the right-side muscles according to SENIAM guidelines [33]: anterior deltoid (AD), middle deltoid (MD), posterior deltoid (PD), bicep brachii (BB), pectoralis major (PM), trapezius (TR), latissimus dorsi (LD), erector spinae (ES). Additionally, an IMU was placed on the right wrist to measure the acceleration. Then, the maximum voluntary contraction (MVC) of each muscle was measured. The order of the conditions and tasks was randomized across participants to avoid order effects. The NASA Task Load Index (NASA-TLX) [34] and Borg CR-10 questionnaires [35] were collected after each trial, to assess the perceived workload (PW) and RPE, respectively. After all trials wearing the exoskeleton ended, the participants filled out the System Usability Scale (SUS) [36]. Fig. 12 presents the experimental procedure for session 2.
@@ 3.5 Data and Statistical Analysis
In session 1, shoulder horizontal and sagittal F/E angles were obtained through a standardized data processing pipeline in VICON. Mean values of the ROM were calculated among all participants in NO HIT-POSE and HIT-POSE conditions. In session 2, acceleration signals were resampled to $1$Hz and variances were calculated through a 30ms time sliding window. The start and end times of the task for each trial were determined by a given variance threshold. EMG signals were processed to compute the linear envelopes by setting them to zero-mean, applying band-pass filter (4th, 20-350Hz), full-wave rectification, low-pass filter (6Hz), and normalization. For each trial, the root mean square (RMS) of each muscle was obtained from the linear envelope via a 30ms time sliding window during the phase of task execution.
Statistical analyses were used to assess the effectiveness of the exoskeleton in reducing muscle activations. The Shapiro-Wilk test was employed to check data normality. If the normal distribution was not satisfied, the Wilcoxon signed-rank test was used to perform statistical analysis, otherwise, the t-test was utilized. Statistical significance was concluded when p $<$ 0.05.